Magical Research in Lords 32

Ryan Powers

August 26, 2001

1 Research Basics

Research is how new magical abilities are learned and designed. Research can be led by a leader, or it can be done without the assistance of a leader. There are four main types of research, these will be covered below. In most cases MFP must be expended to perform the research. The exception to this is if a leader is performing the research the MFP do not need to be spent (but even then it would be a good idea.) Research also improves the nations Magical Tech Level.

To perform research not the research being conducted, and how many MFP are being spent in the notes section of your orders sheet. IF a leader is participating, put the research in the leaders orders and in the notes section.

Research works like investments. Even failed research is not a total loss. Further attempts in the same research will receive benefits from prior attempts.

2 Types of research

When researching, one of the following types must be chosen.

2.1 General

In general research, there is no specific goal. This provides the largest boost to the nations Magical Tech Level. It also provides a moderate chance of learning an ability at random. Useful for learning abilities no one has yet seen, though this can be done other ways as well.

2.2 Question

This form a research allows a question regarding the way magic works to be asked. If successful, an answer will be provided. The more specific the question the higher the success rate will be. Since this also moves the theoretical knowledge of magic forward, it provides a substantial boost to the Magical Tech Level, though not quite as much as the above. It can also provide a new ability relating to the question, but this is rare.

2.3 Class

In this form of research a class of abilities is researched. Once could research combat magics for example, or new leader spells relating to diplomacy. If successful a new ability will be learned, almost always in the class researched. If there is no ability in that class and success is achieved, the option to create one may be given to the player. This created ability will of course be subject to approval and alteration by the GM. This ability provides a small boost to the Magical Tech Level.

2.4 Specific

This is used to learn an ability for which the name is already known. It can be used to learn any of the spells or operations presented in the rules as well as any others whose name is known. This means you should guard the names of your abilities closely, since if someone else knows its name, they may more easily learn it. Specific research provides a high rate of success, but provides the smallest boost to the nation’s Magical Tech Level.

3 Leaders and Research

A leader may be put in charge of a research project. This action requires at least 5 AP, more can be spent if desired. This is the only case when no MFP need be spent, it is assumed in that case that the leader is doing the research on their own. This will reduce the chances of success considerably. It is when the resources of the nation (MFP) are combined under a leader with a high Magic Rating that the most spectacular results can be achieved. Especially if much time is also invested on the part of the leader.